# Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding

### v1

@@ V-F Effectiveness on More LVLMs
V-G Divergence Between Images
@@
We observe that the similarity between images exhibits no strong correlation with performance. Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements. As discussed in Sec.IV-A, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships. This information exhibits limited overlap across real-world images. We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.VI. Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs. In practical applications, the random selection approach offers greater convenience, substantially enhancing the usability of our method.
In Tab. VI, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency. We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion. Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
Figure 6: The correlation between similarity and performance. We conduct this experiment on the validation set of CHAIR. Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
@@
V-H Threshold of Detrimental Priors
@@
VI Conclusion
In this paper, we focus on the language priors issue in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic characteristics of the model and remain consistent across different images. Based on this observation, we propose Cross-Images Contrastive Decoding (CICD)—a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures validate the effectiveness of CICD, particularly in image captioning tasks where language priors are most pronounced.
The simplicity and efficiency of CICD give it substantial practical potential, enabling synergistic integration with other hallucination mitigation methods to further enhance LVLM performance. This constitutes our primary direction for future work.

### v2

@@ C.1 Effectiveness on Larger LVLMs
Table 8: This experiment was conducted by LLaVA-1.5. Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively. w/o Image represents the contrastive distribution that is constructed by removing the input image.
Figure 6: The correlation between similarity and performance. We conducted this experiment on the validation set of CHAIR. Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
C.2 Divergence Between Images
@@
We observe that the similarity between images exhibits no strong correlation with performance. Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements. As discussed in Sec.4.1, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships. This information exhibits limited overlap across real-world images. We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.8. Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs. In practical applications, the random selection approach offers greater convenience, substantially enhancing the usability of our method.
In Tab. 8, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency. We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion. Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
@@
C.3 Threshold of Detrimental Priors